OpenStudy (anonymous):

Am I right? 2(x+6)-10=12+4(x-1) 2x+12-10=12+4x-4 2x+4x=12+4+12-10 6x = 18 6/18 x=3

OpenStudy (anonymous):

no, when you take 4x to right, it becomes -4x

OpenStudy (anonymous):

Then, it's -2x=18?

OpenStudy (anonymous):

2x+12-10=12+4x-4 2x-4x=12-4-12+10 -2x=6 x=-3

OpenStudy (anonymous):

Thanks you.... c: